Project

General

Profile

2016-01-08-Operations

Operations report

  • DC and MOMs
    • Victoria, Paul Hodgson, Henry and Craig are now DCs (last 3 have gone through STFC safety training, Victoria will do so early February)
    • Working well together and with me and the MOMs.
    • DC roster is now on the Operations page on micemine
    • MOMs are no longer needed 24/7 around the year. The post will be rearranged around the user cycles, and will be slightly shorter (3 weeks instead of 4)
    • MOM/DC handover procedure defined and tested
  • ISIS Schedule :
    • 2015/04 : Feb 16 - Mar 25
    • 2016/01 : Apr 12 - May 20
    • 2016/02 : June 28 - July 27
    • 2016/03 : Sep 15 - Oct 28
    • 2016/04 : Nov 15 - Dec 16
  • 2015/04
    • Plan to take straight track data with LiH in March. There will be little running in Feb, apart from expert-led startup runs.
    • Planning to run 16 hours / 24
      • no absorber data (both for baseline comparison with LiH, and alignment)
      • LiH straight track data
    • Need to call for shifters in the next week or so
  • 2016/01
    • Run Plan for April depends on whether the spectrometer solenoids will be operational
    • no SSU or SSD : Data-taking may not be necessary
    • no SSD :
      • channel matching and optics validation
      • beam-based alignment
      • input emittance measurements
      • scattering measurements
    • both SSU and SSD :
      • channel matching and optics validation
      • beam-based alignment
      • input emittance measurements
      • scattering measurements
      • emittance reduction measurements
  • 2016/02 - 2016/04
    • depends on how much channel commissioning work we have done in April. Still being discussed.